Find the solution of Non-Exact D.E





Given:





(x+4y^3)dy-ydx=0

1
Expert's answer
2021-12-09T14:04:58-0500

Let us find the general solution of the differential equation "(x + 4y^3)dy - ydx =0" which is equivalent to "- yx' +x + 4y^3=0." 

Note that "y=0" is a solution. Let us divide both part of the last differential equation by "-y^2."Then we get "\frac{1}{y}x' -\frac{1}{y^2}x =4y." 

It follows that "(\frac{1}{y}x)' = 4y," and hence "\frac{1}{y}x=2y^2+C." 

We conclude that the general solution is "x=2y^3+Cy" or "y=0."